thomas is 1/4 old as his father . If the sum of their two ages is 60, then how old is thomas
Answer by edjones(8007)

Let f=father's age.
f + f/4 = 60
4f+f=240 multiply each side by 4.
5f=240
f=48
48/4=12 Thomas' age.